You've taken a $7300 loan for college at a rate of 7% compounded semiannually. How much do you owe on the loan after 3 years?

Answer by jorel1380(3719)

After 3 years, your loan would be:
A=7300*(1+ .07/2)^6
So:
ln A=ln 7300 + ln (1+0.035)^6=ln 7300+6 ln(1.035)
ln A=8.89562963 + 0.20640856
ln A=9.1020381874404768739922614157774
A=e^9.1020381874404768739922614157774=$8973.56388 as the amount owed after 3 years. ☺☺☺☺